# From Cavitation to Astrophysics: Explicit Solution of the Spherical Collapse Equation

### v3

@@ I Introduction
[…]
@@
Physical system Schematic Meaning of $R$ Time scale $T_{0}$
$\gamma$
$\tau$
@@
Cavitation bubble Spherical cavity of constant (possibly zero) internal pressure, imploding under the pressure of an incompressible inviscid liquid without surface tension
Bubble radius
$R_{0}\sqrt{\rho/\Delta p}$, where $\rho$ is the density of the liquid and $\Delta p=p_{\infty}-p_{v}$, with far-field liquid pressure $p_{\infty}$ and constant bubble pressure $p_{v}$
$-4$
$\approx 0.91468$
Generalisation to a spherical bubble in $N\geq 3$ dimensions [e.g., 14]
$(N-2)^{-1/2}R_{0}\sqrt{\rho/\Delta p}$
$\!\!\!\!\!-N-1$
 (Eq. 16)
@@
Gravitational collapse Collapse of a uniform sphere, also known as spherical top-hat, subjected only to gravitational forces
Radius of sphere
$\sqrt{3/(4\pi G\rho_{0})}$, where $\rho_{0}$ is the initial density at the onset of the collapse and $G$ is the gravitational constant
$-2$
$\pi/\sqrt{8}\approx 1.11072$
@@
Two-body collision Two point masses in gravitational free-fall toward each other, on a straight line without angular momentum
Distance between the centres of the two masses
$R_{0}^{3/2}/\sqrt{G(M_{1}+M_{2})}$, where $M_{i}$ denote the two masses and $G$ is the gravitational constant
$-2$
$\pi/\sqrt{8}\approx 1.11072$
@@
Free-fall in spherical potential Radial free-fall in a spherically symmetric gravitational power law potential, $\phi\sim R^{\alpha}$, with index $\alpha\neq 0$
Distance from centre of potential
$\sqrt{(L/R_{0})^{\alpha}/\alpha}\cdot R_{0}/V$, where $L$ and $V$ are the length and velocity scales of the potential $\phi(R)={\rm sgn}(\alpha)\cdot V^{2}\cdot(R/L)^{\alpha}$
$\alpha-1$
 (Eq. 16)
@@
Same for a logarithmic potential, e.g., the potential of a singular isothermal sphere
$R_{0}/V$, where $V$ is the velocity scale in $\phi(R)=V^{2}\ln(R/L)$
$-1$
$\sqrt{\pi/2}\approx 1.25331$
@@
Free-fall in uniform field Straight fall in a constant acceleration field, such as a vertical drag-free drop of an object on Earth
Height from the ground
$\sqrt{R_{0}/g}$, where $g$ is the norm of the constant acceleration, e.g., $g=9.81\rm m\,s^{-2}$ on Earth
00
$\sqrt{2}\approx 1.41421$
@@
Relativistic orbital decay Orbital decay by gravitational radiation of two masses at non-relativistic velocities, initially forced on circular orbits
Distance between the centres of the two masses
$\frac{5c^{5}R_{0}^{4}}{64\sqrt{3}G^{3}M_{1}M_{2}(M_{1}+M_{2})}$, with $M_{i}$ the two masses, $G$ the gravitational constant and $c$ the speed of light
$-7$
$\approx 0.74683$
@@
Harmonic oscillator Motion of a mass attached to an ideal spring, initially at rest in an extended or compressed state
Distance from the equilibrium position
$\sqrt{M/K}$, where $M$ is the mass and $K$ the spring constant
$+1$
$\pi/2\approx 1.57080$
@@
Polytrope Density profile of a self-gravitating polytropic gas in one dimension in hydrostatic equilibrium
$\rho^{1/n}$, where $\rho$ is the gas density and $n$ the index of $p=K\rho^{1+1/n}$
$\sqrt{(1+n)KR_{0}^{1-n}/(4\pi G)}$ is a length scale set by the parameters of the polytropic equation of state $p=K\rho^{1+1/n}$
$n$
(Eq. 16)
@@
Table 1: Textbook examples of systems governed by Equations (1) and (2). Moving objects described by $R(T)$ are shown in blue, whereas static objects described by this equation are shown in red. $T_{0}$ is the characteristic physical scale (normally a characteristic time), such that Equation (1) reduces to the dimensionless Equation (5) upon normalisation via Equation (3). The collapse time is $T_{c}$, satisfying $R(T_{c})=0$, is given by $T_{0}\tau$. Where $T_{0}$ depends on $R_{0}$, the latter is the maximum, initial value of $R$. In the case of the polytrope (bottom row), $R$ and $T$ have different meanings: $R:=\rho^{1/n}$ is a measure of the gas density and $T$ is the distance from the centre of mass. For most examples, analogous cases with other forces, such as electric forces instead of gravitational ones, can be found.
@@ II.1 Dimensionless form
It is understood that dots above $r(t)$ now denote derivatives with respect to $t$, rather than $T$. To help the following derivations, we introduce the variable $\tau:=T_{c}/T_{0}$ as a shorthand for the dimensionless collapse time.
@@ III.1.1 Cases $\gamma\neq-1$
This case can be solved by reduction to quadrature, a standard method expressing the solution of ODEs in terms of integrals. For general second order ODEs, $\ddot{f}(t)=f(\dot{r},r,t)$, the integrals found in this way must be evaluated numerically. Interestingly, however, the limited form of Equation (5) allows us, through adequate substitutions, to express the emerging integrals through the incomplete beta function $B(x;\alpha,\beta)$ (derivation in Appendix A). This special function benefits from an extensive mathematical literature available for further analytical manipulations of the solution. Moreover, by introducing the shorthands
@@ III.2 Explicit solution $r(t)$
The significance of Equation (20b) relies in the fact that the regularised incomplete beta function $I(x;\alpha,\beta)$ is the cumulative density function of the beta distribution, a common distribution function in probability theory, related to, but not to be confused with, the beta function. Hence, the function $Q(x;\alpha,\beta)$ is the quantile function of the beta distribution. This connection to the beta distribution, which has been intensively studied in probability theory, makes Equation (20b) a practical solution for further analytical manipulations. Moreover, the quantile function $Q(x;\alpha,\beta)$ is readily accessible to all modern programming languages, making this solution easy to implement. Two examples in Python and R are given in Appendix B.
@@ III.3 Noteworthy special solutions
To my knowledge, Equations (21da–c) are the only explicit closed-form solutions, i.e., solutions in terms of commonly accepted basic functions. These three functions respectively describe an arc of a circle (Equation 21a), a parabola (Equation 21b), and a harmonic oscillation (Equation 21c). Textbook examples for the latter two are given in Table 1.
Equation (21d) is a special case of the undamped and unforced Duffing equation [15], for which general solutions in terms of the Jacobi elliptic family have recently been derived [16].
@@ IV.4 General symmetry
The symmetry of Equation (25) can help identify new solutions based on existing ones, with the most trivial example being the derivation of Equation (21a) from Equation (21b), and vice versa.
@@ Acknowledgements.
I thank Dr. Martin Bruderer, Dr. Aaron Ludlow, Dr. Zachari Slepian, Dr. Dmitry Sinelshchikov and Prof. Stefani Mancas for insightful suggestions that have inspired and improved this work. I am also grateful to Dr. Roberto Iacono for bringing to my attention the convenience of using the regularised incomplete beta function in transitioning from Equation (15) to Equation (20a), bypassing an earlier method invoking the hypergeometric function. I also acknowledge an anonymous referee of an earlier draft, for pointing me toward the substitutions discussed in Section LABEL:sss:subsymmetry. Last but not least, I extend my gratitude to Dr. Mohamed Farhat and Prof. Andrea Prosperetti for many past discussions, which were a source of inspiration in writing this paper. I am a recipient of an Australian Research Council Future Fellowship (FT190100083) funded by the Australian Government.

### v4

@@ I Introduction
It may be surprising that the 21st century still holds critically important differential equations, which admit compact, but barely known explicit solutions. In this paper, I discuss such a family of equations that govern all physical systems described by a one-dimensional time-dependent coordinate $R(T)$, subject to a restoring force that varies as a power law of that coordinate. Formally, this is expressed by the non-linear ordinary differential equation (ODE)
@@
Physical system Schematic Meaning of $R$ Time scale $T_{0}$ $\gamma$ $\tau$
@@
Cavitation bubble Spherical cavity–empty or with low constant inner pressure–imploding under the pressure of an incompressible inviscid liquid without surface tension [Uncaptioned image] Bubble radius $R_{0}\sqrt{\rho/\Delta p}$, where $\rho$ is the density of the liquid and $\Delta p=p_{\infty}-p_{v}$, with far-field pressure $p_{\infty}$ and constant bubble pressure $p_{v}\ll p_{\infty}$ $-4$ $\approx 0.91468$
Generalisation to a spherical bubble in $N\geq 3$ dimensions [e.g., 14] $(N-2)^{-1/2}R_{0}\sqrt{\rho/\Delta p}$ $\!\!\!\!\!-N-1$  (Eq. 17)
@@
Gravitational collapse Collapse of a uniform sphere, also known as spherical top-hat, subjected only to gravitational forces [Uncaptioned image] Radius of sphere $\sqrt{3/(4\pi G\rho_{0})}$, where $\rho_{0}$ is the initial density at the onset of the collapse and $G$ is the gravitational constant $-2$ $\pi/\sqrt{8}\approx 1.11072$
@@
Two-body collision Two point masses in gravitational free-fall toward each other, on a straight line without angular momentum [Uncaptioned image] Distance between the centres of the two masses $R_{0}^{3/2}/\sqrt{G(M_{1}+M_{2})}$, where $M_{i}$ denote the two masses and $G$ is the gravitational constant $-2$ $\pi/\sqrt{8}\approx 1.11072$
@@
Free-fall in spherical potential Radial free-fall in a spherically symmetric gravitational power law potential, $\phi\sim R^{\alpha}$, with index $\alpha\neq 0$ [Uncaptioned image] Distance from centre of potential $\sqrt{(L/R_{0})^{\alpha}/\alpha}\cdot R_{0}/V$, where $L$ and $V$ are the length and velocity scales of the potential $\phi(R)={\rm sgn}(\alpha)\cdot V^{2}\cdot(R/L)^{\alpha}$ $\alpha-1$  (Eq. 17)
@@
Same for a logarithmic potential, e.g., the potential of a singular isothermal sphere $R_{0}/V$, where $V$ is the velocity scale in $\phi(R)=V^{2}\ln(R/L)$ $-1$ $\sqrt{\pi/2}\approx 1.25331$
@@
Free-fall in uniform field Straight fall in a constant acceleration field, such as a vertical drag-free drop of an object on Earth [Uncaptioned image] Height from the ground $\sqrt{R_{0}/g}$, where $g$ is the norm of the constant acceleration, e.g., $g=9.81\rm m\,s^{-2}$ on Earth 00 $\sqrt{2}\approx 1.41421$
Acceleration by a dipole charge Straight fall of a free charge toward a fixed dipole of charges, separated by a distance $\varepsilon\ll R$ along the trajectory [Uncaptioned image] Distance between free charge and dipole centre $\sqrt{MR_{0}^{4}/(4\varepsilon k_{e}Qq)}$, where $M$ is the mass of the free charge $Q\!>\!0$, $q\!>\!0$ the dipole charge and $k_{e}$ the Coulomb constant $-3$ $1$
@@
Relativistic orbital decay Orbital decay by gravitational radiation of two masses at non-relativistic velocities, initially on circular orbits [Uncaptioned image] Distance between the centres of the two masses $\frac{5c^{5}R_{0}^{4}}{64\sqrt{3}G^{3}M_{1}M_{2}(M_{1}+M_{2})}$, with $M_{i}$ the two masses, $G$ the gravitational constant and $c$ the speed of light $-7$ $\approx 0.74683$
@@
Harmonic oscillator Motion of a mass attached to an ideal spring, initially at rest in an extended or compressed state [Uncaptioned image] Distance from the equilibrium position $\sqrt{M/K}$, where $M$ is the mass and $K$ the spring constant $+1$ $\pi/2\approx 1.57080$
@@
Polytrope Density profile of a self-gravitating polytropic gas in one dimension in hydrostatic equilibrium [Uncaptioned image] $\rho^{1/n}$, where $\rho$ is the gas density and $n$ the index of $p=K\rho^{1+1/n}$ $\sqrt{(1+n)KR_{0}^{1-n}/(4\pi G)}$ is a length scale set by the parameters of the equation of state $p=K\rho^{1+1/n}$ $n$ (Eq. 17)
@@
Table 1: Textbook examples of systems governed by Equations (1) and (2). Moving objects described by $R(T)$ are shown in blue, whereas static objects described by this equation are shown in red. $T_{0}$ is the characteristic physical scale (normally a characteristic time), such that Equation (1) reduces to the dimensionless Equation (5) upon normalisation via Equation (3). Where $T_{0}$ depends on $R_{0}$, the latter is the maximum, initial value of $R$. In the case of the polytrope (bottom row), $R$ and $T$ have different meanings: $R:=\rho^{1/n}$ is a measure of the gas density and $T$ is the distance from the centre of mass. The last column shows the dimensionless collapse time, $\tau$, such that the physical collapse time is $T_{c}=\tau T_{0}$, where $R(T_{c})=0$. For most examples, analogous cases with other forces, such as electric forces instead of gravitational ones, can be found.
@@ II.1 Dimensionless form
It is understood that dots above $r(t)$ now denote derivatives with respect to $t$, rather than $T$. To help the following derivations, we introduce the variable
$\tau:=\frac{T_{c}}{T_{0}}$ (7)
as a shorthand for the dimensionless collapse time.
@@ III.1.1 Cases $\gamma\neq-1$
This case can be solved, starting from the integral of motion in Equation (9a), by reduction to quadrature, a standard method expressing the solution of ODEs in terms of integrals. For general first order ODEs, $\dot{r}(t)=f(r(t),t)$, the integrals found in this way must be evaluated numerically. Interestingly, however, the limited form of Equation (5) allows us, through adequate substitutions, to express the emerging integrals through the incomplete beta function $B(x;\alpha,\beta)$ (derivation in Appendix A). This special function benefits from an extensive mathematical literature available for further analytical manipulations of the solution. Moreover, by introducing the shorthands
@@ III.2 Explicit solution $r(t)$
The significance of Equation (21b) relies in the fact that the regularised incomplete beta function $I(x;\alpha,\beta)$ is the cumulative density function of the beta distribution, a common distribution function in probability theory, related to, but not to be confused with, the beta function. Hence, the function $Q(x;\alpha,\beta)$ is the quantile function of the beta distribution. This connection to the beta distribution, which has been intensively studied in probability theory, makes Equation (21b) a practical solution for further analytical manipulations. Moreover, the quantile function $Q(x;\alpha,\beta)$ is readily accessible to all modern programming languages, making this solution easy to implement. Two examples in Python and R code are given in Appendix B.
@@ III.3 Noteworthy special solutions
To my knowledge, Equations (22da–c) are the only explicit closed-form solutions, i.e., solutions in terms of commonly accepted basic functions. These three functions respectively describe an arc of a circle (Equation 22a), a parabola (Equation 22b), and a harmonic oscillation (Equation 22c). Textbook examples of corresponding physical systems are given in Table 1.
Equation (22d) is a special case of the Duffing equation [15], describing an undamped and unforced anharmonic oscillator. General solutions in terms of the Jacobi elliptic family have recently been derived [16].
@@ IV.4 General symmetry
The symmetry of Equation (26) can help identify new solutions based on existing ones, an example being the derivation of Equation (22a) from Equation (22b), and vice versa.
@@ Acknowledgements.
I thank Dr. Martin Bruderer, Dr. Aaron Ludlow, Dr. Zachari Slepian, Dr. Dmitry Sinelshchikov, Dr. Roberto Iacono and Prof. Stefani Mancas for insightful suggestions that have inspired and improved this work. I also acknowledge both anonymous referees for their comments and positivity, as well as an anonymous referee of an earlier draft, for pointing out the substitutions discussed in Section IV.4. Last but not least, I extend my gratitude to Dr. Mohamed Farhat and Prof. Andrea Prosperetti for many past discussions, which were a source of inspiration in writing this paper. I am a recipient of an Australian Research Council Future Fellowship (FT190100083) funded by the Australian Government.